OpenStudy (anonymous):

will fan and medal Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8) and Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (-8, 0), (1, 5) if you can

OpenStudy (aripotta):

first you have to find the slope between those two points, (7,-3) and (4,-8) slope formula:\[m~=\frac{ (y _{2}-y _{1}) }{ (x _{2}-x _{1}) }\]

OpenStudy (aripotta):

then, choose one of the two points, and plug it into point-slope formula, along with the slope you just found. point-slope formula: \[y-y _{1}=m(x-x _{1})\] where m is the slope, y1 is the y-coordinate of the point, and x1 is the x-coordinate of the same point

OpenStudy (aripotta):

lastly, convert your point-slope equation into standard form.